# Quantum Universality in Composite Systems: A Trichotomy of Clifford Resources

### v1

@@ Abstract
The efficient classical simulation of Clifford circuits constitutes a fundamental barrier to quantum advantage, typically overcome by injecting explicit non-Clifford “magic” resources. We demonstrate that for high-dimensional quantum systems (qudits), the resources required to break this barrier are strictly governed by the number-theoretic structure of the Hilbert space dimension $d$. By analyzing the adjoint action of the Clifford group, we establish a classification of single-qudit universality as a trichotomy. (I) For prime dimensions, the Clifford group is a maximal finite subgroup, and universality is robustly achieved by any non-Clifford gate. (II) For prime-power dimensions, the group structure fragments, requiring tailored diagonal non-Clifford gates to restore irreducibility. (III) Most notably, for composite dimensions with coprime factors, we demonstrate that standard entangling operations alone—specifically, generalized intra-qudit CNOT gates—generate the necessary non-Clifford resources to guarantee a dense subgroup of $\mathrm{SU}(d)$ without explicit diagonal magic injection. Our proofs rely on a new geometric criterion establishing that a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element with projective distance strictly less than $1/2$ from the identity. These results suggest that “coprime architectures”—hybrid registers combining incommensurate dimensions—can bypass the standard overhead of magic state distillation.
@@ I Introduction
What makes a quantum computer more powerful than a classical one? The standard narrative points to interference and entanglement, yet the Gottesman-Knill theorem reveals that highly-entangled states can be efficiently simulated classically when generated by Clifford operations alone [Gottesman1998, AaronsonGottesman2004]. The missing resource that lifts computation beyond classical reach is typically identified as “magic”: non-stabilizer states or non-Clifford gates that enable quantum universality—the ability to approximate any unitary operator to arbitrary precision by composing elements from a finite gate set [Kitaev1997, Deutsch1995]. This paper reveals that for quantum systems of composite dimension magic can emerge from the interplay of classical operations, with the local Hilbert space dimension unfolding as a previously unrecognized computational resource.
Consider a quantum register composed of subregisters with coprime local dimensions—say, a qutrit ($d=3$) coupled to a ququart ($d=4$). Coupling between them can be attained via the generalized intra-qudit controlled-addition gate: a permutation matrix that adds the control value to the target modulo $d$. Notably, this operation is classical and specified entirely by discrete modular arithmetic. Yet we prove that when such intra-qudit gates connect subregisters of coprime dimension, the group they generate together with local Clifford operations is dense in the special unitary group. Universality emerges from arithmetic alone.
This inverts the conventional understanding of quantum resources. In standard qubit architectures, the single-qubit Clifford group supplemented by CNOT gates remains finite—it is the multi-qubit Clifford group, efficiently simulable by the Gottesman-Knill theorem. Achieving universality demands a dedicated non-Clifford resource, such as the $T$-gate or magic states prepared through costly distillation protocols [Bravyi2005]. Our results show that this overhead is an artifact of homogeneous architectures built from identical subsystems, not an intrinsic feature of quantum computation. When local dimensions are coprime, the “free” entangling operations that any architecture provides are already sufficient.
The phenomenon has a transparent algebraic origin. The generalized CNOT is a classical reversible gate. However, when connecting subsystems of coprime dimensions $p$ and $q$, it acts on local Pauli operators to produce diagonal gates. By Bézout’s identity, these gates construct fractional phases $2\pi/pq$—inaccessible to either local Clifford group. Hence, the generation of dense subgroups follows from gates that individually preserve the computational basis. This suggests viewing local dimension as a computational resource. Unlike prime-power dimensions like $d=4$, coprime dimensions like $d=6$ allow the “mixing” of incompatible modular arithmetics, creating quantum resources from the arithmetic structure alone.
A finitely generated subgroup of $\mathrm{SU}(d)$ is dense if and only if it is infinite and acts irreducibly on the Lie algebra $\mathfrak{su}(d)$ via the adjoint representation. When applied to extensions of the Clifford group, this criterion identifies the specific obstructions to universality. In composite dimensions, unlike the prime case, the Clifford group is not maximal; therefore, a non-Clifford gate may generate a strictly larger yet still finite group. To rule out this possibility and ensure infiniteness, we introduce a geometric criterion: a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element within projective distance $1/2$ of the identity.
@@
Related work. The structure of universality for qubit systems is well understood: the Clifford group supplemented by any non-Clifford gate—typically the $T$-gate from the third level of the Clifford hierarchy—generates a dense subgroup of $\mathrm{SU}(2)$ [Gottesman1999, Nebe2001]. This “Clifford + TT” paradigm underlies modern fault-tolerant quantum computation, where magic states enable non-Clifford operations through gate teleportation [Bravyi2005, Knill2004]. For higher-dimensional systems, the situation is less complete. Universality criteria for qudits of prime dimension have been established [Howard2012, Campbell2012], and the maximality of the Clifford group in these cases follows from classical results on finite linear groups [Nebe2001]. However, composite dimensions with mixed prime factors have not been systematically studied, even as experimental platforms increasingly access non-binary levels [RingbauerEtAl2022, Low2025, Chicco_2023].
@@
Universality for arbitrary dimension $d$, not based on the Clifford group, was established through both constructive and algebraic frameworks. Previous work proved universality by decomposing elements of $\mathrm{SU}(d^{n})$ into two-level unitaries, which were further expressed through phase and controlled operations [LuoWang2014]. Alternative constructions employed the generators of $\mathfrak{su}(d)$, where universal gate sets were generated via exponentiation of string operators, to produce $\mathrm{U}(d^{n})$ [VlasovJMP2002, VlasovSPIE2003].
Main results. We classify the resources required for single-qudit universality by the prime factorization of $d$:
@@
    Prime dimensions $\bm{(d=p).}$ The Clifford group acts irreducibly on the Lie algebra and is a maximal finite subgroup of $\mathrm{PSU}(d)$. We prove maximality holds if and only if $d$ is prime: adding any non-Clifford gate forces universality. In this regime, every non-Clifford gate is automatically magic.
@@
    Prime power dimensions $\bm{(d=p^{m},\,m\geq 2).}$ The adjoint representation becomes reducible. The Lie algebra fragments into invariant subspaces indexed by divisibility. Universality requires specific diagonal non-Clifford gates—members of the $T_{s}$ family whose phases restore irreducibility. Not every non-Clifford gate suffices; the arithmetic obstruction must be overcome by deliberate design.
Cases I and II share a common strategy: universality via diagonal non-Clifford gates (the $T_{s}$ family).
@@
    Composite dimensions with coprime factors $\bm{(d=d_{1}\cdots d_{n}).}$ Universality emerges from intra-qudit gates alone. When $d$ has pairwise coprime factors, generalized CNOTs connecting subsystems construct diagonal gates via Bézout’s identity. These gates satisfy both density conditions: they restore irreducibility by mixing the local Clifford structures, and they certify infiniteness by lying within projective distance $1/2$ of the identity. No explicit magic injection is required. In contrast to prime dimensions, the failure of maximality allows for intermediate finite groups, meaning that a gate can be non-Clifford without being magic.
Case III is our main contribution. In prime and prime-power dimensions (Case I and II), universality requires the explicit injection of diagonal non-Clifford gates ($T_{s}$). In contrast, coprime dimensions achieve universality using only generalized CNOTs. Although these are classical reversible permutations, they are non-Clifford resources. The arithmetic mismatch allows these permutation gates to generate a dense subgroup, making universality an intrinsic feature of the hardware that does not require explicit diagonal magic.
Structure of the paper. Section II establishes the algebraic framework: the density criterion, the structure of the generalized Pauli and Clifford groups, and the decomposition of composite-dimensional systems via the Chinese Remainder Theorem. Section III develops the geometric criterion for infiniteness and its spectral characterization. Sections IV, V, and VI analyze the three regimes of the trichotomy in detail. Section VII discusses implications of our findings and open problems.
@@ II.1 Criteria for Universality
The focus of this work is primarily on establishing single-qudit universality. This focus is justified by the fact that the complexity of the universality problem resides almost entirely in the local structure. Once the single-qudit case has been completely characterized, extending this to universal quantum computation on an array of qudits requires only a single additional gate $V$ that breaks the preservation of the local tensor structure (an entangling resource). This reduction is formalized by the following seminal result:
Theorem 2 ([Brylinski2002]).
@@
Consequently, establishing universality reduces to identifying when a finite set of gates generates a dense subgroup of the special unitary group. Proving density directly via topological arguments is often hard. The fundamental difficulty lies in distinguishing whether a finitely generated subgroup is dense or merely trapped within a proper, infinite closed Lie subgroup of lower dimension.
@@
for $g\in\mathrm{SU}(d)$ and $X\in\mathfrak{su}(d)$. Since unitary conjugation preserves the Hilbert-Schmidt inner product, this action is unitary. A crucial property is that the irreducibility of this action is insensitive to complexification: the action on $\mathfrak{su}(d)$ is irreducible if and only if the complexified action on $\mathfrak{sl}(d,\mathbb{C})$ is irreducible.
@@
This lemma leads to one criteria for single-qudit universality.
@@ II.2 The Clifford Group and its Structure
The Heisenberg group $H(d)$ (sometimes called the generalized Pauli group or just the Pauli group) consists of:
@@
The adjoint action of any representative $U\in C(d)$ maps Pauli operators to Pauli operators, inducing an automorphism on the index space. This action corresponds to the special linear group over the ring $\mathbb{Z}/d\mathbb{Z}$. This structural rigidity is characterized by the short exact sequence [Tolar_2018]:
@@
An important property for our analysis of composite systems is that this structure factorizes over coprime dimensions, allowing us to treat prime-power sectors independently.
@@
Recall that by the Chinese Remainder Theorem (CRT), the map $\chi:\mathbb{Z}/d\mathbb{Z}\to\mathbb{Z}/{d_{1}}\mathbb{Z}\times\mathbb{Z}/{d_{2}}\mathbb{Z}$ given by $x\mapsto(x\bmod d_{1},x\bmod d_{2})$ is a ring isomorphism.
@@
    Hadamard ($\bm{H}$): The Fourier transform kernel $\omega^{xy}$ factorizes because the inner product in $\mathbb{Z}_{d}$ decomposes into the sum of local inner products (modulo integer scaling), yielding $H_{d}\cong H_{d_{1}}\otimes H_{d_{2}}$.
@@
    Phase Gate ($\bm{P}$): The gate applies a phase $\omega^{Q(x)}$, where $Q(x)=x(x+\rho_{d})/2$ is a quadratic polynomial. Since $\chi$ is a ring homomorphism, it preserves polynomial evaluation: computing $Q(x)$ modulo $d$ is equivalent to computing pairs $(Q(x_{1})\bmod d_{1},Q(x_{2})\bmod d_{2})$. Consequently, the diagonal operator factorizes as $P_{d}\cong P_{d_{1}}\otimes P_{d_{2}}$.
@@
III A Geometric Criterion for Infiniteness

### v2

@@ Abstract
The Clifford group is efficiently classically simulable, and universality is obtained by supplementing it with non-Clifford resources. We determine which single-qudit gates suffice to achieve universality. We show that the structure of such resources is governed by the prime factorization of the qudit dimension $d$. Using the adjoint action on the space of complex trace-zero matrices, we relate density to irreducibility together with an infiniteness criterion, yielding a trichotomy based on the factorization of $d$. When $d$ is prime, any non-Clifford gate generates a dense subgroup of the determinant-one unitaries. If $d$ is a prime power, the adjoint action is reducible, and universality requires gates that couple the resulting invariant subspaces. For composite $d$ with pairwise coprime factors, generalized intra-qudit controlled-NOT gates connecting the factors already suffice. These findings suggest that “composite architectures”—hybrid registers combining incommensurate dimensions—offer a route to bypass the standard overhead associated with magic-state injection.
@@ I Introduction
A central question in quantum information science is understanding the computational advantage of quantum computers over classical architectures. While quantum entanglement is often cited as the source of such an advantage, the Gottesman-Knill theorem reveals a fundamental limitation: highly-entangled states generated by Clifford operations alone can be efficiently simulated classically [19, 1]. This result extends to higher-dimensional qudits, where efficient classical simulation of Clifford circuits has been established [15, 12]. The missing resource that lifts computation beyond classical reach is typically identified as magic: non-stabilizer states or non-Clifford gates that enable quantum universality—the ability to approximate any unitary operator to arbitrary precision by finite gate compositions [24]. Establishing conditions under which the Clifford group in high-dimensional systems yields universal quantum computation remains an open problem. In this work, we analyze this question for the single-qudit Clifford group.
[…]
Establishing universality amounts to proving that the generated group is dense in $\mathrm{SU}(d)$, the group of unitary operators with determinant one. We address this topological problem via the adjoint representation of the group on $\mathfrak{sl}(d,\mathbb{C})$. A finitely generated subgroup of $\mathrm{SU}(d)$ is dense if and only if it is infinite and acts irreducibly on the Lie algebra $\mathfrak{sl}(d,\mathbb{C})$ (Theorem 4). The extent to which the Clifford group satisfies these conditions is governed by the prime factorization $d=p_{1}^{m_{1}}\cdots p_{k}^{m_{k}}$ (Theorem 24 and Theorem 30). This factorization dictates the algebraic obstruction, establishing a natural trichotomy.
When $d=p$ is prime, the Clifford group exhibits maximal rigidity. We prove that its adjoint action is irreducible if and only if $d$ is prime (Theorem 24). Furthermore, the group is projectively maximal if and only if $d$ is prime: any strictly larger finite subgroup must differ only by scalar phases (Theorem 25). Consequently, adding any non-Clifford gate forces universality (Corollary 27). However, when the factorization includes higher powers ($m_{i}\geq 2$) or multiple distinct primes ($k>1$), this rigidity breaks down. The adjoint action becomes reducible, fragmenting the Lie algebra into invariant subspaces indexed by divisibility properties of $d$ (Theorem 30). Achieving universality, therefore, requires resources capable of coupling these subspaces to restore irreducibility.
[…]
We establish density by verifying two conditions. First, irreducibility requires characterizing the invariant subspaces of $\mathfrak{sl}(d,\mathbb{C})$ under the Clifford group’s adjoint action. We classify these subspaces via the symplectic orbit structure of $\mathrm{SL}(2,\mathbb{Z}/d\mathbb{Z})$ and demonstrate that our resource gates mix the resulting components. Second, infiniteness is established through a simple criterion: a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element within projective distance $1/2$ of the identity. This criterion reduces the verification to a computation over the eigenvalues.
@@
Related work. The structure of universality for qubit systems is well understood: the Clifford group supplemented by any non-Clifford gate—typically the $T$-gate from the third level of the Clifford hierarchy [18]—generates a dense subgroup of $\mathrm{SU}(2)$ [20, 34]. This “Clifford + TT” paradigm underlies modern fault-tolerant quantum computation, where magic states enable non-Clifford operations through gate teleportation [7, 25].
For higher-dimensional systems, the situation is less complete. Universality criteria for qudits of prime dimension have been established [23, 9], and the maximality of the Clifford group in these cases follows from classical results on finite linear groups [34]. However, prime-power and composite dimensions have not been systematically studied, even as experimental platforms increasingly access non-binary levels [35, 31, 10].
@@
Universality for arbitrary dimension $d$, not based on the Clifford group, has been established through both constructive and algebraic frameworks. Previous work proved universality by decomposing elements of $\mathrm{SU}(d^{n})$ into two-level unitaries, which were further expressed through phase and controlled operations [32]. Alternative constructions employed the generators of $\mathfrak{su}(d)$, where universal gate sets were generated via exponentiation of string operators, to produce $\mathrm{U}(d^{n})$ [39, 40].
Main results. We provide a complete characterization of the resources required for single-qudit universality, based on the prime factorization of the local dimension $d$. The classification reflects a hierarchy of algebraic obstructions within the Clifford group:
@@
    Prime dimensions $\bm{(d=p).}$ No obstruction. The Clifford group acts irreducibly on the Lie algebra and is a maximal finite subgroup of $\mathrm{SU}(d)$ modulo scalar phases. We prove that both maximality and irreducibility of the Clifford group hold if and only if $d$ is prime: adding any non-Clifford gate forces universality. In this case, every non-Clifford gate necessarily has magic.
@@
    Prime-power dimensions $\bm{(d=p^{m},\,m\geq 2).}$ Reducibility obstruction. The adjoint representation becomes reducible, decomposing the Lie algebra into invariant subspaces indexed by divisibility. The Clifford group acts transitively within each subspace but cannot connect them. Universality requires a non-Clifford gate that restores irreducibility by coupling such subspaces. Two classes of gates accomplish this:
    • Diagonal gates from the $T_{s}$ family with $s\nmid d$.
    • Non-affine permutations, such as the transposition (0 1)(0\;1).
    Not every non-Clifford gate suffices; one must choose a gate with the correct algebraic properties.
@@
    Composite dimensions with coprime factors $\bm{(d=d_{1}\cdots d_{k}).}$ Obstruction resolved by arithmetic. The Clifford group decomposes as a direct product $\mathcal{C}_{d}\cong\mathcal{C}_{d_{1}}\times\cdots\times\mathcal{C}_{d_{k}}$ of the local Clifford groups, and consequently fails to act irreducibly on the global Lie algebra (Proposition 8). However, generalized intra-qudit CNOTs connecting coprime factors automatically restore irreducibility and construct the necessary diagonal gates (Proposition 47). When $k=2$, a single CNOT coupling the two factors suffices; for $k>2$, multiple CNOTs connecting adjacent factors are required. These gates satisfy both density conditions: they restore irreducibility by coupling the local Clifford structures, and they certify infiniteness by lying within projective distance $1/2$ of the identity (Theorem 48). No explicit diagonal magic injection is required. Again, not every non-Clifford gate suffices (see conditions in Theorem 42).
The trichotomy demonstrates that universality requirements depend strictly on how $d$ factorizes into prime powers. For prime dimensions, the Clifford group is maximally rigid—any non-Clifford gate suffices. For prime powers, reducibility demands gates that restore irreducibility. For coprime dimensions, intra-qudit CNOTs connecting the factors restore irreducibility and provide the necessary non-Clifford structure. In all three cases, the prime factorization determines both the obstruction to universality and the algebraic properties required to overcome it.
Structure of the paper. Section II establishes our algebraic framework, including the density criterion based on the adjoint representation, the structure of the Clifford group via its symplectic quotient, and the factorization of composite systems. Section III introduces an infiniteness criterion: a subgroup with irreducible adjoint action is infinite if it contains a non-scalar element within projective distance $1/2$ of the identity. Sections IVVI systematically establish universality for each case of the trichotomy. Section VII concludes with implications and open questions.
@@ II.1 Criterion for Universality
The focus of this work is primarily on establishing single-qudit universality. This focus is justified by the fact that the complexity of the universality problem resides almost entirely in the local structure. Once the single-qudit case has been completely characterized, extending this to universal quantum computation on an array of qudits requires only a single additional gate $V$ that breaks the preservation of the local tensor structure (an entangling resource). This reduction is formalized by the following result:
Theorem 2 ([8]).
@@
Consequently, establishing universality reduces to identifying when a finite set of gates generates a dense subgroup of the special unitary group. Proving density directly via topological arguments is often hard. The main difficulty lies in distinguishing whether a finitely generated subgroup is dense or trapped within a proper, infinite closed Lie subgroup of lower dimension.
@@
for $g\in\mathrm{SU}(d)$ and $X\in\mathfrak{su}(d)$. Since unitary conjugation preserves the Hilbert-Schmidt inner product, this action is unitary. An important property is that the irreducibility of this action is insensitive to complexification: the action on $\mathfrak{su}(d)$ is irreducible if and only if the complexified action on $\mathfrak{sl}(d,\mathbb{C})$ is irreducible.
@@
This lemma leads to the following criterion for single-qudit universality.
@@ II.2 The Clifford Group and its Structure
The Heisenberg group $H(d)$ (sometimes called the generalized Pauli group or just the Pauli group [15, 12]) consists of:
@@
The adjoint action of any representative $U\in C(d)$ maps Pauli operators to Pauli operators, inducing an automorphism on the index space. This action corresponds to the special linear group over the ring $\mathbb{Z}/d\mathbb{Z}$. This structural rigidity is characterized by the short exact sequence [37]:
@@
The finite Clifford group factorizes over coprime dimensions—a property previously identified by Looi and Griffiths [29] for stabilizer states—which allows us to reduce the problem to independent prime-power sectors.
@@
Recall that by the Chinese Remainder Theorem, the map $\chi:\mathbb{Z}/d\mathbb{Z}\to\mathbb{Z}/{d_{1}}\mathbb{Z}\times\mathbb{Z}/{d_{2}}\mathbb{Z}$ given by $x\mapsto(x\bmod d_{1},x\bmod d_{2})$ is a ring isomorphism.
@@
    Hadamard ($H$): The Fourier transform kernel $\omega^{xy}$ factorizes because the inner product in $\mathbb{Z}_{d}$ decomposes into the sum of local inner products (modulo integer scaling), yielding $H_{d}\cong H_{d_{1}}\otimes H_{d_{2}}$.
@@
    Phase Gate ($P$): The gate applies a phase $\omega^{Q(x)}$, where $Q(x)=x(x+\rho_{d})/2$ is a quadratic polynomial. Since $\chi$ is a ring homomorphism, it preserves polynomial evaluation: computing $Q(x)$ modulo $d$ is equivalent to computing pairs $(Q(x_{1})\bmod d_{1},Q(x_{2})\bmod d_{2})$. Consequently, the diagonal operator factorizes as $P_{d}\cong P_{d_{1}}\otimes P_{d_{2}}$.
@@
II.3 Permutations in the Clifford Group
While the Clifford group acts mainly by creating superpositions via the Hadamard gate, it also contains a subgroup of classical reversible operations corresponding to permutations of the computational basis. Characterizing this subgroup is essential for identifying which classical gates constitute non-Clifford resources.
Let $S_{d}$ be the symmetric group of degree $d$. For any permutation $\pi\in S_{d}$, we define the associated permutation matrix $P_{\pi}\in\mathrm{U}(d)$ by its action on the basis states: $P_{\pi}|x\rangle=|\pi(x)\rangle$.
Proposition 9.
A permutation matrix $P_{\pi}$ belongs to the Clifford group $\mathcal{C}_{d}$ if and only if the underlying permutation $\pi:\mathbb{Z}_{d}\to\mathbb{Z}_{d}$ is an invertible affine transformation over the ring $\mathbb{Z}_{d}$. That is, $\pi$ is of the form:
$\pi(x)=ax+b\pmod{d}$, (8)
where $a\in\mathbb{Z}_{d}^{\times}$ (a unit in $\mathbb{Z}_{d}$) and $b\in\mathbb{Z}_{d}$. Consequently, the subgroup of Clifford permutations is isomorphic to the affine group $\mathrm{Aff}(1,\mathbb{Z}_{d})$.
Proof.
$(\Rightarrow)$ Assume $P_{\pi}\in\mathcal{C}_{d}$. By definition, conjugation by $P_{\pi}$ must map the Heisenberg group $H(d)$ to itself.
Consider the action on the Pauli-$Z$ operator, which is diagonal with entries $\omega^{x}$ where $\omega=e^{2\pi i/d}$. Conjugation permutes the diagonal entries:
$P_{\pi}ZP_{\pi}^{\dagger}=\sum_{x\in\mathbb{Z}_{d}}\omega^{x}|\pi(x)\rangle\langle\pi(x)|=\sum_{y\in\mathbb{Z}_{d}}\omega^{\pi^{-1}(y)}|y\rangle\langle y|.$
For the resulting diagonal operator to belong to $H(d)$, it must be of the form $\omega^{c_{0}}Z^{k}$ for some $k,c_{0}\in\mathbb{Z}_{d}$. Comparing the diagonal entries, we require $\omega^{\pi^{-1}(y)}=\omega^{c_{0}+ky}$ for all $y\in\mathbb{Z}_{d}$. This implies:
$\pi^{-1}(y)\equiv ky+c_{0}\pmod{d}$.
Since $\pi^{-1}$ is a bijection, the coefficient $k$ must be a unit in $\mathbb{Z}_{d}$; otherwise, the map $y\mapsto ky+c_{0}$ would fail to be injective. Thus, $\pi^{-1}$ is an affine map with invertible linear part, and consequently $\pi$ itself must be affine: $\pi(x)=ax+b$ with $a\in\mathbb{Z}_{d}^{\times}$.
$(\Leftarrow)$ Conversely, assume $\pi(x)=ax+b$ with $a\in\mathbb{Z}_{d}^{\times}$. We verify the conjugation action on the generators $Z$ and $X$ of the Heisenberg group.
For the Pauli-$Z$ operator, a direct calculation using $\pi^{-1}(y)=a^{-1}(y-b)$ gives:
$P_{\pi}ZP_{\pi}^{\dagger}=\sum_{y\in\mathbb{Z}_{d}}\omega^{a^{-1}(y-b)}|y\rangle\langle y|=\omega^{-a^{-1}b}Z^{a^{-1}},$
which belongs to $H(d)$.
For the operator $X$, we compute:
$P_{\pi}XP_{\pi}^{\dagger}|y\rangle$ $=P_{\pi}X|\pi^{-1}(y)\rangle=P_{\pi}|\pi^{-1}(y)+1\rangle$
$=|\pi(\pi^{-1}(y)+1)\rangle.$
Substituting $\pi^{-1}(y)=a^{-1}(y-b)$, the argument of $\pi$ becomes:
$\pi\big(a^{-1}(y-b)+1\big)=a\cdot a^{-1}(y-b)+a+b=y+a.$
Thus, $P_{\pi}XP_{\pi}^{\dagger}=X^{a}$. Since $a$ is a unit, $X^{a}$ generates the same cyclic subgroup as $X$ and belongs to $H(d)$.
Since $P_{\pi}$ maps the generators of $H(d)$ to elements of $H(d)$, we conclude $P_{\pi}\in\mathcal{C}_{d}$. ∎
Remark 10.
The order of the affine group is $|\mathrm{Aff}(1,\mathbb{Z}_{d})|=\varphi(d)\cdot d$, where $\varphi$ denotes Euler’s totient function.
For $d=2$, we have $|S_{2}|=2$ and $|\mathrm{Aff}(1,\mathbb{Z}_{2})|=1\cdot 2=2$. For $d=3$, we have $|S_{3}|=6$ and $|\mathrm{Aff}(1,\mathbb{Z}_{3})|=2\cdot 3=6$. In both cases, the groups coincide: every permutation is affine and thus a Clifford operation.
However, for $d\geq 4$, the inequality $\varphi(d)\cdot d<d!$ implies the strict inclusion $\mathrm{Aff}(1,\mathbb{Z}_{d})\subsetneq S_{d}$. Consequently, non-affine permutations exist. For instance, in $d=4$, the transposition (0 1)(0\;1)—which swaps $|0\rangle$ and $|1\rangle$ while fixing $|2\rangle$ and $|3\rangle$—is a non-Clifford gate despite being a classical reversible operation.
II.4 Diagonal Gates and the $T_{s}$-gate Family
To characterize diagonal non-Clifford operators, for any dimension $d$, we introduce a generic family of diagonal gates.
Definition 11.
Definition 12.
$\delta_{\zeta}(x,y)=\frac{\zeta(x+y)}{\zeta(x)\zeta(y)}.$ (10)
Lemma 13.
Proof.
The conjugation action on the shift operator $X^{y}$ is given by:
We specifically focus on the following $T_{s}$ family.
Definition 14 (The $T_{s}$-gate Family).
For instance, $T_{d}$ recovers the Pauli $Z$ gate (order $d$), and $T_{2d}$ corresponds to the generalized phase gate. To determine the “Cliffordness” of $T_{s}$, we analyze its conjugation action on $X$:
For $T_{s}$ to belong to the Clifford group $\mathcal{C}_{d}$, it must normalize the Heisenberg group. Since the result is proportional to $X$, this condition holds if and only if the phase $e^{2\pi i/s}I$ is in the Heisenberg group:
  • Odd Dimensions: The phases are generated by $\omega=e^{2\pi i/d}$. Thus, $T_{s}\in\mathcal{C}_{d}\iff s\mid d$.
  • Even Dimensions: The Clifford group generates phases of order $2d$. Thus, $T_{s}\in\mathcal{C}_{d}\iff s\mid 2d$.
This leads to a unified criterion for when $T_{s}$ constitutes a non-Clifford resource.
Proposition 15.
III A Criterion for Infiniteness